OpenStudy (anonymous):

Specify the collection where f is continuous f(x) = x^3-3x+2

OpenStudy (anonymous):

\[(-\infty,\infty)\]

OpenStudy (anonymous):

I got some other ex of that kind, what methode do I have to use the 2e is: f(x)= (x-2)/(x+3)

OpenStudy (anonymous):

hold it. how did the polynomial become a rational function?

OpenStudy (anonymous):

That is another excercise.

OpenStudy (anonymous):

could you recheck the function.

OpenStudy (anonymous):

Yes, we are seeing raional functions.

OpenStudy (anonymous):

no. but what you have here is: \[f(x)=x^3-3x+2\]

OpenStudy (anonymous):